# MENTOR: Guiding Hierarchical Reinforcement Learning with Human Feedback and Dynamic Distance Constraint

### v1

@@ 4.1 Framework formalization
RLHF uses pairwise comparison to train a reward model which can be used as the reward function for the high-level policy. The training process consists of 1) extracting pairs $(s_{1},sg_{1},g)$ and $(s_{2},sg_{2},g)$ from the high-level replay buffer. 2) Human annotators provide pairwise feedback $v$ which 0 prefers $(s_{1},sg_{1},g)$ , and 1 prefers $(s_{2},sg_{2},g)$. Preference can be assessed by the distance to the environment goal $g$. 3) After collecting the feedback dataset into the reward model buffer $B_{hf}$, we can train the reward model $r_{hf}$ by optimizing a modified Bradley-Terry objective [Bradley and Terry, 1952]. We define the possibility that human prefers $(s_{1},sg_{1},g)$ to $(s_{2},sg_{2},g)$:
$P[(s_{1},sg_{1})\succ(s_{2},sg_{2})|g]$ (1)
$\quad=\frac{\exp(r_{hf}(s_{1},sg_{1},g))}{\sum_{i=1}^{2}\exp(r_{hf}(s_{i},sg_{i},g))}.$
Then the modified Bradley-Terry objective is as follows:
$\max_{r_{hf}}\mathbb{E}_{(s_{1},sg_{2},s_{1},sg_{2},g,v)\sim B_{hf}}$ (2)
$\quad[v\log(P[(s_{1},sg_{1})\succ(s_{2},sg_{2})|g])+$
$\quad(1-v)\log(P[(s_{2},sg_{2})\succ(s_{1},sg_{1})|g])].$
In optimizing the high-level policy, the high-level reward function is set to be $r_{hi}=r_{hf}(s,sg,g)$. Nevertheless, two issues arise with this direct implementation of $r_{hf}$. (1) Training the reward model with all off-policy data is inefficient. When the low-level policy is already capable of achieving subgoals, comparing these subgoals is unnecessary for reward model training. For the issue, we propose training the reward model with near-policy data, using $(s,sg,g)$ pairs sampled from recent episodes. This approach focuses on the most relevant data, reducing computational and model complexity. (2) In some cases, the high-level policy may fall into a local optimum due to the lack of a pretraining step to improve the reward function by comparing more diverse samples. Therefore, it is important to encourage the high-level to explore. We introduced some exploration techniques such as RND and maximum entropy (MaxEnt) [Mei et al., 2020]. We apply MaxEnt and set $r_{hi}=r_{hf}(s,sg,g)-\beta\cdot\mathbb{E}_{sg\sim\pi_{hi}(\cdot|s,g)}\log\pi_{hi}(sg|s,g)$. $\beta$ is a small constant.
Low-level Policy. Sparse rewards hinder the low-level ability to quickly achieve subgoals. Current research employs hindsight relabelling, which limits exploration capabilities and leads to repeated exploration. Introducing RND can mitigate repeated exploration by promoting novel discoveries, though its direct implementation may result in instability. RND’s exploration incentives might lead to neglecting the task of subgoal completion. To address this issue, we propose EED, a dual policy program, consisting of an exploration policy $\pi_{lo}^{e}$ and a base policy $\pi_{lo}$, both sharing the same data buffer. During interactions with the environment, we employ the exploration policy $\pi_{lo}^{e}$ and store the experiences in the shared replay buffer. For policy updates, both the exploration and base policies undergo a hindsight relabelling process for relabeled transition data. Then, the exploration policy is updated using reward $r_{g}(s^{\prime},sg)+f_{rnd}(s^{\prime})$. The base policy is updated using the reward $r_{g}(s^{\prime},sg)$. This method allows the base policy to learn from novel exploration data without conducting the exploration task, concentrating on achieving the intended subgoals.

### v2

@@ III-A1 Subgoal generation using RLHF
RLHF uses pairwise comparison to train a reward model which can be used as the reward function for the high-level policy. The training process consists of (1) extracting pairs $(s_{1},g^{\text{sub}}_{1},g^{\text{env}})$ and $(s_{2},g^{\text{sub}}_{2},g^{\text{env}})$ from the high-level replay buffer. (2) Human annotators provide pairwise feedback $v$ which 0 prefers $(s_{1},g^{\text{sub}}_{1},g^{\text{env}})$ , $1$ prefers $(s_{2},g^{\text{sub}}_{2},g^{\text{env}})$ and $0.5$ for same preference. Preference can be assessed by the distance to the environment goal $g^{\text{env}}$. (3) After collecting the feedback dataset into the reward model buffer $B_{\text{hf}}$, we can train the reward model $r_{\text{hf}}$ by optimizing a modified Bradley-Terry objective [43]. We define the possibility that human prefers $(s_{1},g^{\text{sub}}_{1},g^{\text{env}})$ to $(s_{2},g^{\text{sub}}_{2},g^{\text{env}})$:
$P\left[(s_{1},g^{\text{sub}}_{1})\succ(s_{2},g^{\text{sub}}_{2})|g^{\text{env}}\right]$ (4)
$\quad=\frac{\exp(r_{\text{hf}}(s_{1},g^{\text{sub}}_{1},g^{\text{env}}))}{\sum_{i=1}^{2}\exp(r_{\text{hf}}(s_{i},g^{\text{sub}}_{i},g^{\text{env}}))}.$
Then, we define two probabilities:
$p_{1}=P[(s_{1},g^{\text{sub}}_{1})\succ(s_{2},g^{\text{sub}}_{2})|g^{\text{env}}],$
$p_{2}=P[(s_{2},g^{\text{sub}}_{2})\succ(s_{1},g^{\text{sub}}_{1})|g^{\text{env}}],$
and the modified Bradley-Terry objective is as follows:
$\max_{r_{\text{hf}}}\mathbb{E}_{(s_{1},g^{\text{sub}}_{1},s_{2},g^{\text{sub}}_{2},g^{\text{env}},v)\sim B_{\text{hf}}}\left[(1-v)\log(p_{1})+v\log(p_{2})\right].$ (5)
In optimizing the high-level policy, the high-level reward function is set to be $r_{\mathrm{hi}}=r_{\text{hf}}(s,g^{\text{sub}},g^{\text{env}})$. In the dynamic process of execution, humans and algorithms engage in ongoing interactions, where high-level policies can be subject to real-time guidance from human operators.
III-A2 Dynamic Distance Constraint for Subgoal Difficulty Adjustment
while humans can provide direction for high-level goals by expressing preferences, they may struggle to define subgoals that align with the low-level policy’s capabilities. It is quite possible that influenced by their cognition, humans might prefer subgoals that are close to the ultimate goal. Consequently, if the reward model learned by human preferences is designed to assign the highest rewards when the subgoal coincides with the goal, it could lead to a scenario where the high-level policy, without considering the low-level’s capacity and basing its decisions solely on human feedback, generates subgoals that rapidly converge towards the goal. This approach could render the subgoals excessively challenging and diminish their instructional value. Thus multi-level simultaneous learning may be uncoordinated.
To solve this issue above, we introduce DDC, whose function is illustrated in Figure 2. This method limits the range of subgoals based on their distance from the current state (the variation in green shading depicted in the figure). Under the function of DDC, the learning process of our framework MENTOR is similar to curriculum learning[44, 45, 46]. Curriculum learning facilitates a sequential learning process where simpler subgoals are mastered before progressing to more complex ones, thus laying a solid foundation for advanced subgoal acquisition. This is achieved through a specific formulation:
$\max_{\pi_{\mathrm{hi}}}$ $\ \mathbb{E}_{g^{\text{sub}}\sim\pi_{\mathrm{hi}}}\left[\sum_{i=0}^{n}r_{\mathrm{hi}}(s_{t_{i}},g^{\text{sub}}_{i},g^{\text{env}})\right]$ (6)
s.t. $\mathcal{H}(g^{\text{sub}}_{i},g^{\text{ach}}_{t_{i}},k):=\max\{d(g^{\text{sub}}_{i},g^{\text{ach}}_{t_{i}})-k,0\}\leq 0,$
where $d(g^{\text{sub}}_{i},g^{\text{ach}}_{t_{i}}):\mathcal{G}\times\mathcal{G}\rightarrow[0,1]$ represents the distance between the subgoal $g^{\text{sub}}_{i}$ and the achieved goal $g^{\text{ach}}_{t_{i}}$. The parameter $k$ sets the subgoal space range. The constraint ensures that the distance between the high-level subgoal and the current achieved goal remains within a range of $k$ lengths, and we can adjust $k$ to progressively reduce the difficulty. However, in scenarios like the Four Rooms domain, the Euclidean distance as $d(\cdot)$ may not accurately assess the difficulty of completing a task. For instance, a goal in the top right might be hard to reach despite a low Euclidean distance, indicating the need for a learned step distance measure. DDL [47] offers a methodology for training distance models by randomly sampling state pairs from trajectories, to approximate the number of steps between states. Nevertheless, it’s crucial to evaluate the step distance between unreachable subgoals and the current state. If the low-level policy fails to reach a subgoal, assigning a high distance value to these unreached subgoals relative to the current state is essential. Neglecting to do so, high-level policies might incorrectly view challenging subgoals as easy, leading to unrealistic subgoal proposals. To tackle this issue, we recommend incorporating extra samples containing unreached subgoals into the distance model objective:
$\min_{d}\frac{1}{2}\mathbb{E}_{\tau\sim\rho_{\pi_{\mathrm{lo}}},i\in[0,L],j\in[i,L],g^{\text{sub}}\sim\pi_{\mathrm{hi}}(\cdot|s_{0},g^{\text{env}})}$ (7)
$\left[(d(s_{i},s_{j})-\frac{|i-j|}{L})^{2}+\mathbbm{1}(s_{L}\neq g^{\text{sub}})(d(s_{i},g^{\text{sub}})-1)^{2}\right].$
The objective is formulated by minimizing the expected loss across trajectories $\tau$, sampled using the low-level policy $\rho_{\pi_{\mathrm{lo}}}$ from recent episodes, and the goals $g^{\text{env}}$ drawn from the environment’s distribution $\rho_{g^{\text{env}}}$. $L$ is the length of the trajectory.
In Equation 6, optimizing is challenging because of the strict constraint. To overcome this difficulty, we utilize the penalty function method [48, 49], which allows us to establish an unconstrained optimization objective:
$\max_{\pi_{\mathrm{hi}}}\mathbb{E}_{g^{\text{sub}}\sim\pi_{\mathrm{hi}}}\sum_{i=0}^{n}[r_{\mathrm{hi}}(s_{t_{i}},g^{\text{sub}}_{i},g^{\text{env}})-\alpha\mathcal{H}(g^{\text{sub}}_{i},g^{\text{ach}}_{t_{i}},k)],$ (8)
where $\alpha$ is a balancing coefficient to adjust the influence of reward from human guidance and distance constraint. $\mathcal{H}(g^{\text{sub}}_{i},g^{ach}_{t_{i}},k)$ serves as a penalty function that imposes a cost when the subgoal $g^{\text{sub}}_{i}$ deviates from the achievable range defined by the parameter $k$. This parameter $k$ acts as a threshold, defining the maximum allowable distance for a subgoal to be attainable. This mechanism ensures that the subgoals chosen by the high-level policy are both in harmony with human guidance and within the operational capacity of the low-level policy.
As the capabilities of the low-level improve, it becomes necessary to dynamically adjust the parameter $k$ to ensure that the difficulty of the subgoals remains appropriately challenging. The adjustments in $k$ lead to different optimization goals for the model, which can cause several issues: (1) The coefficient needs to strike a balance between being large enough to ensure constraints are met and small enough to avoid excessive punishment. If the penalty coefficient is set too high, it may overly restrict the solution search space, causing the algorithm to miss the optimal solution or fail to converge. Conversely, if the penalty coefficient is too low, it may not effectively enforce the constraints, leading to solutions that do not satisfy the actual problem constraints. Therefore, a static coefficient may not preserve the optimal balance between human guidance and constraints when $k$ varies. (2) Different values of $k$ lead to distinct distributions of subgoals. Sampling data pairs solely from the comprehensive offline dataset does not ensure that the reward model accurately assigns rewards to subgoals under the current $k$. (3) Assuming policy convergence based on a specific $k$, it’s challenging to quickly re-optimize the high-level policy when switching to a new $k$. In response to these challenges, we have implemented three designs to ensure stability in policy updates when there are changes in $k$.
Automatic balancing coefficient. We implement a dynamic balancing coefficient which will be updated in real-time to maintain a balance between the rewards and distance constraint.
To effectively incorporate the adjustment into MENTOR, we need to optimize two parameters simultaneously: high-level policy $\theta_{\mathrm{hi}}$ and balancing coefficient $\alpha$, converting our maximization into a dual problem:
$\min_{\alpha\geq 0}\max_{\pi_{\mathrm{hi}}}\mathbb{E}_{g^{\text{sub}}\sim\pi_{\mathrm{hi}}}$ $\sum_{i=0}^{n}[r_{\mathrm{hi}}(s_{t_{i}},g^{\text{sub}}_{i},g^{\text{env}})-$ (9)
$\alpha\mathcal{H}(g^{\text{sub}}_{i},g^{\text{ach}}_{t_{i}},k)].$
The distance constraint function $\mathcal{H}(\cdot)$ guarantees that the subgoals within a distance $k$. This function treats subgoals that are within the distance $k$ as having zero effects. However, in the case of $\alpha$, we want it to decrease when this constraint does not work, and increase when the high-level policies make decisions that do not satisfy this constraint. Since $\mathcal{H}(\cdot)$ is always greater than 0 and the update gradient for $\alpha$ is singular, we need to eliminate the $max(\cdot)$ function from $\mathcal{H}(\cdot)$ to achieve automatic updates for $\alpha$. We update the policy $\pi_{\mathrm{hi}}$ firstly and then coefficient $\alpha$ following modification for the optimization:
$\pi_{\mathrm{hi}}^{*}=\arg\max_{\pi_{\mathrm{hi}}}\mathbb{E}_{g^{\text{sub}}\sim\pi_{\mathrm{hi}}}$ $\sum_{i}^{n}[r_{\mathrm{hi}}(s_{t_{i}},g^{\text{sub}}_{i},g^{\text{env}})-$ (10)
$\alpha\mathcal{H}(g^{\text{sub}}_{i},g^{\text{ach}}_{t_{i}},k)],$
$\alpha^{*}$ $=\arg\min_{\alpha}\mathbb{E}_{g^{\text{sub}}\sim\pi_{\mathrm{hi}}}\sum_{i}^{n}\left[-\alpha\left(d(g^{\text{sub}}_{i},g^{\text{ach}}_{t_{i}})-k\right)\right].$ (11)
Near-policy sampling. The probability distribution of the data varies depending on the sampling policy and the different values of $k$, even though all the data exist in an offline experience pool. It is inefficient to train the reward model using all off-policy data. Once the low-level policy has become proficient at achieving subgoals, it becomes redundant to use these subgoals for the training of the reward model, which is intended to enhance the current policy. To address this issue, we propose a new method that involves training the reward model with near-policy data, using $(s,g^{\text{sub}},g)$ pairs that are sampled from recent episodes. This approach allows the reward model to focus on the data that is most relevant to the current policy, enabling a more accurate evaluation of the current policy’s performance. Moreover, training a reward model that can accurately evaluate the behavior of the current policy requires fewer samples and training iterations, which results in reduced computational consumption.
Near-policy sampling can also be efficiently executed by merely preserving data from the newest episodes or their respective indices within the experience pool, and subsequently, randomly extracting from these recent datasets when gathering the comparison samples for the reward function training.
High-level exploratory mechanism. As depicted in Figure 2, the high-level policy may encounter a local optimum after proposing a subgoal in the lower-left region (left panel) and the low-level policy has adapted to achieve it, particularly upon adjusting the $k$ parameter. This policy is guided by a reward model trained on data from the replay buffer and human feedback, lacking initial pre-training. With incremental $k$ values, the high-level policy, potentially already converged, may limit its exploration to a less diverse dataset. This restricted exploration could prevent the identification and storage of superior subgoals in the replay buffer, thereby impeding the enhancement of the reward model and leading to policy training stagnation. To counteract this, it is crucial to promote high-level exploration. We have integrated exploration techniques such as RND and Maximum Entropy (MaxEnt), as detailed in [50]. Specifically, we have adopted the MaxEnt approach, defining the high-level reward as $r_{\mathrm{hi}}=r_{\text{hf}}(s,g^{\text{sub}},g^{\text{env}})-\beta\cdot\log\pi_{\mathrm{hi}}(g^{\text{sub}}|s,g^{\text{env}})$, where $\beta$ is a small constant.
III-B Exploration-Exploitation Decoupling in Low-level Policy
Although the high-level can provide easier goals for the low-level through RLHF, these goals also serve as guidance for the low-level to move forward. However, sparse rewards hinder the low-level ability to quickly achieve subgoals. RLHF is well-suited for the high-level in HRL, but it is not applicable to the low-level. If the low-level policy incorporated RLHF, the noise and uncertainty in reward model could hinder the completion of the task. Existing technologies may use hindsight relabeling, but before the low-level policy explores the subgoals, this technology cannot guarantee that the agent learns how to complete the subtasks. Due to the limited exploration ability of the low-level policy, the agent may fall into local optima and repeatedly explore meaningless areas. Introducing RND can mitigate repeated exploration by promoting novel discoveries, though its direct implementation may result in instability. RND’s exploration incentives might lead to neglecting the task of subgoal completion.
To address the issue, we propose EED, a dual-policy program, consisting of an exploration policy $\pi_{\mathrm{lo}}^{e}$ and a base policy $\pi_{\mathrm{lo}}$, both sharing the same data buffer. During interactions with the environment, we employ the exploration policy $\pi_{\mathrm{lo}}^{e}$ and store the experiences in the shared replay buffer. For policy updates, both the exploration and base policies undergo a hindsight relabelling process for relabeled transition data.
Subsequently, the exploratory policy is refined by optimizing the following objective function:
$\max_{\pi_{\mathrm{lo}}^{e}}$ $\mathbb{E}_{\tau\sim\pi_{\mathrm{lo}}^{e}}\left[\sum_{i=1}^{L}r_{\text{g}}(s_{t},g^{\text{sub}})+r_{\text{intri}}(s_{t})\right].$ (12)
This formulation incorporates an additional element, $r_{\text{intri}}(s_{t})$, which introduces curiosity into the reward structure, thereby fostering exploration. Conversely, the base policy is updated by optimizing the following objective function:
$\max_{\pi_{\mathrm{lo}}}\mathbb{E}_{\tau\sim\pi_{\mathrm{lo}}^{e}}\left[\sum_{i=1}^{L}r_{\text{g}}(s_{t},g^{\text{sub}})\right].$ (13)
This approach enables the base policy to assimilate novel insights from the exploratory data without the inherent burden of the exploration process itself, thus enabling it to focus on the attainment of the defined subgoals.
III-C MENTOR Process
Overall, this paper introduces an innovative HRL algorithm, which proficiently identifies an approach for configuring the reward function, RLHF for high-level and EED for low-level. Furthermore, the paper addresses the challenge of inter-layer coordination in HRL by proposing a novel optimization with dynamic distance constraint. In detail, our framework MENTOR works as the following pseudo-code in Algorithm 1. During the interaction with the environment (from line 4 to line 17), a high-level policy $\pi_{\mathrm{hi}}$ is responsible for selecting a subgoal $g^{\text{sub}}$ (line 6). Once the subgoal is determined, a low-level exploration policy, denoted as $\pi_{\mathrm{lo}}^{e}$, is utilized to execute actions until the subgoal is successfully achieved. This process of selecting subgoals and executing actions continues until the end of the episode. The data obtained from the interaction with the environment, both at the high-level and low-level, is stored in two replay buffers known as $B_{\mathrm{hi}}$ and $B_{\mathrm{lo}}$. The model update process is implemented in lines 19 to 23. The high-level policy in Algorithm 3 updates $\pi_{\mathrm{hi}}$ and $\alpha$ alternately. As for the low-level policy, it involves updating the RND model, applying hindsight, updating the low-level base policy $\pi_{\mathrm{lo}}$, adding the RND bonus in the batch data, and updating the exploration policy $\pi_{\mathrm{lo}}^{e}$. From lines 21 to 22, preference tuple pairs are sampled from the data of the last few episodes in $B_{\mathrm{hi}}$, and human labels are obtained and stored in $B_{\text{hf}}$. These batches are then used to update $r_{\text{hf}}$ and rewrite the reward data in the high-level buffer $B_{\mathrm{hi}}$. Finally, distance training data is sampled from the data of the last few episodes in $B_{\mathrm{hi}}$ and used to update the distance model $d$. From lines 25 to 34, it tests the low-level base policy success rate on the subgoal $g^{\text{sub}}$ given by high-level policy and adjusts the parameters $k$.
Algorithm 1 MENTOR
1:  Input: initial state distribution $\rho(s)$, goal distribution $\rho(g)$, terminal threshold $\epsilon$
2:  Initialize high-level policy $\pi_{\mathrm{hi}}$, low-level base policy $\pi_{\mathrm{lo}}$, low-level exploration policy $\pi_{\mathrm{lo}}^{e}$, reward model $r_{\text{hf}}$, RND model $f_{\text{rnd}}$, distance model $d$, balancing coefficient $\alpha$, constraint hyper-parameters $k$, high-level buffer $B_{\mathrm{hi}}$, preference buffer $B_{\text{hf}}$ and low-level buffer $B_{\mathrm{lo}}$, a high threshold $high\_threshold$ for the test success rate at the value $k$, a low threshold $low\_threshold$ for the test success rate at the value $k$.
3:  for $i=0$ to $M$ do
4:     receive initial state $s_{t}$ and initial $g_{\text{env}}$ from the environment, $s_{t}\leftarrow\rho(s),g_{\text{env}}\leftarrow\rho(g_{\text{env}})$
5:     repeat
6:        select a subgoal $g^{\text{sub}}$leftarrow $\pi_{\mathrm{hi}}(s_{t},g_{\text{env}})$
7:        store high-level state $s_{\mathrm{hi}}\leftarrow s_{t}$
8:        repeat
9:           select action $a_{t}\leftarrow$ $\pi_{\mathrm{lo}}^{e}(s_{t},g^{\text{sub}})$
10:           execute $a_{t}$ and observe next state $s_{t+1}$
11:           compute low-level reward $r_{\mathrm{lo}}=r_{\text{g}}(s_{t+1},g^{\text{sub}})$
12:           store $(s_{t},a_{t},s_{t+1},g^{\text{sub}},r_{l})$ to low-level buffer $B_{\mathrm{lo}}$
13:           $s_{t}\leftarrow s_{t+1}$
14:        until $\mathbbm{1}(\|g^{\text{sub}}-s_{t}\|_{2}<\epsilon)$ or end of episode
15:        calculate the high reward $r_{\mathrm{hi}}\leftarrow r_{\text{hf}}(s_{\mathrm{hi}},g^{\text{sub}},g_{\text{env}})$
16:        append $(s_{\mathrm{hi}},g^{\text{sub}},s_{t},g_{\text{env}},r_{\mathrm{hi}})$ to $B_{\mathrm{hi}}$
17:        $s_{\mathrm{hi}}\leftarrow s_{t}$
18:     until end of episode
19:     TrainHighPolicy($\pi_{\mathrm{hi}}$, $d$, $\alpha$, $B_{\mathrm{hi}}$) Algorithm 2
20:     TrainLowPolicy($\pi_{\mathrm{lo}}$, $\pi_{\mathrm{lo}}^{e}$, $f_{\text{rnd}}$, $B_{\mathrm{lo}}$) Algorithm 3
21:     near-policy query preference tuples pairs from $B_{\mathrm{hi}}$ and store preferences in $B_{\text{hf}}$
22:     update $r_{\text{hf}}$ on batches from $B_{\text{hf}}$ using Equation 5 and rewrite the reward in $B_{\mathrm{hi}}$
23:     near-policy sample distance batches from $B_{\mathrm{lo}}$ and update distance model $d$ using Equation 7
24:     $success\_rate\leftarrow\{\}$
25:     for $j=0$ to $N_{1}$ do
26:        sample goal $g$ from environment and select a subgoal $g^{\text{sub}}$
[…]
28:     end for
29:     if average($success\_rate$) $\geq$ $high\_threshold$ then
30:        increase $k$ value
31:     end if
32:     if average($success\_rate$) $<$ $low\_threshold$ then
33:        decrease $k$ value
34:     end if
35:  end for
Algorithm 2 TrainHighPolicy
1:  Input: high-level base policy $\pi_{\mathrm{hi}}$, distance model $d$, balancing coefficient $\alpha$, high-level buffer $B_{\mathrm{hi}}$
2:  for $i=0$ to $N_{2}$ do
3:     sample a batch $B$ from buffer $B_{\mathrm{hi}}$
4:     use batch $B$ to update policy $\pi_{\mathrm{hi}}$ following Equation 10
5:     use batch $B$ to update dual variable $\alpha$ following Equation 11
6:  end for
Algorithm 3 TrainLowPolicy
1:  Input: low-level base policy $\pi_{\mathrm{lo}}$, low-level exploration policy $\pi_{\mathrm{lo}}^{e}$, RND model $f_{\text{rnd}}$, low-level buffer $B_{\mathrm{lo}}$
2:  for $i=0$ to $N_{3}$ do
3:     sample a batch $B$ from buffer $B_{\mathrm{lo}}$
4:     use batch $B$ to update RND model $f_{\text{rnd}}$ using Equation 1
5:     using hindsight to rewrite the goal and reward using $r_{\text{g}}$ in $B$
6:     use batch $B$ to update policy $\pi_{\mathrm{lo}}$ using Equation 13
7:     add RND bonus $f_{\text{rnd}}$ to reward in $B$ using Equation 3
8:     use batch $B$ to update exploration policy $\pi_{\mathrm{lo}}^{e}$ using Equation 12
9:  end for